OpenStudy (anonymous):

x+2=√(x-4) solve for x.

OpenStudy (anonymous):

square both sides and then simplify, you should end up with a quadratic

OpenStudy (anonymous):

1) Squaring both sides: (x+2)^2 = (x-4) 2) Expanding (x+2)^2 (x^2) + (4x) + 4 = (x-4) 3) Simplifying the expression: (x^2) + (4x-x) + (4+4) = 0 (x^2) + 3x + 8 = 0 4) Solving for x = ?

OpenStudy (anonymous):

so my final answer would be x^2+3x+8 rite?

OpenStudy (anonymous):

Not quite

OpenStudy (anonymous):

wait y not?

OpenStudy (anonymous):

it said solve for x...so find solutions for x that satisfy that equation

OpenStudy (anonymous):

The directions say to solve for x Which should look like this: x=...

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

seems you will get complex roots.

OpenStudy (anonymous):

oh like w/ imaginary #'s?

OpenStudy (anonymous):

yea imaginary number

OpenStudy (anonymous):

try using the quadratic formula

OpenStudy (anonymous):

ugh! but that's not really an option

OpenStudy (anonymous):

ok i'll try

OpenStudy (anonymous):

did you copy the question down correctly?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

it was exactly: x+2=√(4-x)

OpenStudy (anonymous):

4-X not X-4

OpenStudy (anonymous):

yes isn't that what I said??

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

oops sorry

OpenStudy (anonymous):